Find three consecutive integers that solve this:
5x=2y+(7+z)

Answer by oberobic(2304)

x, y, and z are 3 are three consecutive integers.
.
They can be defined as:
x
x+1
x+2
.
BUT, we don't know if that is the x, y, z order or not.
The three integers could be z, y, x. Or z, x, y. Or y, x, z...
.
Assuming x, y, z order:
.
5x = 2y + (7+z)
5x = 2(x+1) + (7 + x +2)
5x = 2x +2 + 7 + x + 2
5x = 3x + 11
2x = 11
x = 11/2
.
That is not an integer.
.
Assuming x=5
5x = 25
2*6 + 7+7 = 26
.
Assuming x=6
5x = 30
2*7 +7+8 =29
.
Are you sure that is the equation?
.
Done.